# MaTe3D: Mask-guided Text-based 3D-aware Portrait Editing

### v3

@@ Abstract
3D-aware portrait editing has a wide range of applications in multiple fields. However, current approaches are limited due that they can only perform mask-guided or text-based editing. Even by fusing the two procedures into a model, the editing quality and stability cannot be ensured. To address this limitation, we propose MaTe3D: mask-guided text-based 3D-aware portrait editing. In this framework, first, we introduce a new SDF-based 3D generator which learns local and global representations with proposed SDF and density consistency losses. This enhances masked-based editing in local areas; second, we present a novel distillation strategy: Conditional Distillation on Geometry and Texture (CDGT). Compared to exiting distillation strategies, it mitigates visual ambiguity and avoids mismatch between texture and geometry, thereby producing stable texture and convincing geometry while editing. Additionally, we create the CatMask-HQ dataset, a large-scale high-resolution cat face annotation for exploration of model generalization and expansion. We perform expensive experiments on both the FFHQ and CatMask-HQ datasets to demonstrate the editing quality and stability of the proposed method. Our method faithfully generates a 3D-aware edited face image based on a modified mask and a text prompt. Our code and models will be publicly released.
Index Terms:
@@
Mask-guided Editing, Text-guided Editing, Diffusion Model, Score Distillation Sampling
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Leveraging the advancements of StyleGAN series [1, 2, 3, 4], EG3D [5] and related works like [6, 7, 8, 9] have successfully created high-quality 3D scenes with consistent views and rich details. Recent 3D-aware image synthesis methods [10, 11, 12] aim to enhance view-consistent image editing capabilities by incorporating additional information like semantic maps. This caters to a range of user needs, including 3D avatar creation, personalized services, and industrial design.
@@
Existing NeRF-based generation methods with editability can be categorized into two classes: mask-guided [10, 11] and text-guided [13, 12]. Both methods rely on a common pipeline: (1) training a 3D-aware generator; (2) exploring various editing methods based on the well-trained generator. Mask-guided methods aim to facilitate user-friendly editing through hand-drawn sketches or semantic label interfaces. For example, IDE-3D [11] enables both high-quality and real-time local face control by manipulating a semantic map. However, current mask-guided approaches primarily emphasize shape manipulation and lack texture control capabilities. On the other hand, text-guided methods strive to produce more expressive results based on textual descriptions by distilling knowledge from large models. LENeRF [12] utilizes CLIP [14] for precise and localized manipulation using text inputs. Nevertheless, existing text-based techniques encounter challenges in accurately controlling shape to achieve desired results(Fig. 6). In this paper, we focus on using both masked-guided and text-guided 3D image editing with high quality and stability. Masks allow for precise spatial modifications, such as removing or altering specific parts of the scene, while text editing allows for semantic changes based on natural language descriptions. Utilizing both mask and text enhances interactivity and control by enabling users to interact with and modify 3D scenes in more intuitive ways.
@@
II Related Works
II-A 3D-aware GAN based Synthesis and Editing
@@ II Related Works
Shocked by NeRF [15] and SIREN [16] in exceeding the past method with regards to multi-view image synthesis, recent methods start to focus on 3D-aware models without multi-view supervision. GRAF [17] and pi-GAN [18] are pioneers in exploring NeRF-based 3D-aware synthesis and render view-consistent results. StyleNeRF [7] and CIPS-3D [8] adopt progressively upsampling to synthesize high-quality 3D-aware images. StyleSDF [6], which merges NeRF with Signed Distance Function (SDF) [19], achieves comparative performance to aforementioned methods and also produces 3D smooth complex shapes. EG3D [5] introduces a hybrid explicit-implicit 3D representation that produces high-resolution image and brings implicitly learned mesh to a new level of meticulosity. However, none of the above methods consider fine-grained locally-controllable face editing with other conditions.
@@
To achieve fine-grained level control with mask, FENeRF [10] pioneers local face editing using decoupled latent codes to generate corresponding facial semantics and texture in a spatially aligned 3D volume with shared geometry. IDE-3D [11] and NeRFFaceEditing [20] allow for semantic mask guided local control of images in real time. CNeRF [21] and 3D-SSGAN [22] aim to achieve part-level editing through the compositional synthesis. For fine-grained editing with text, LENeRF [12] achieves local manipulation via establishing 3D attention map via CLIP [14]. StyleGANFusion [23] and FaceDNeRF [24] distill knowledge from diffusion model to enable manipulate Face NeRFs in zero-shot learning based on text prompts.
II-B Diffusion Models Guided Editing
@@
In recent years, diffusion models have made great strides in the field of text-to-2D generation. Some works have begun to use diffusion model priors to edit 3D scenes through text guidance. For example, Instruct-NeRF2NeRF [25] employs Instruct-Pix2Pix [26] to modify NeRF content using language instructions. The approach optimizes both the underlying scene and iterative optimization, resulting in a 3D scene that conforms to the input text. Blended-NeRF [27] uses a pre-trained language image model to synthesize objects based on textual prompts and blend them into existing scenes using 3D multi-layer perceptron (MLP) models and volume blending techniques for natural appearance.
@@
Score Distillation Sampling (SDS) from DreamFusion [28] provides a new way to apply diffusion models to 3D scenes. DreamEditor [29] allows controllable editing of neural fields using textual prompts through SDS. FocalDreamer [30] merges shapes with editable parts based on SDS for fine-grained editing in desired regions. SKED [31], which edits 3D shapes represented by NeRFs, utilizes two guiding sketches without altering existing fields much, supporting object creation or modification within designated sketch regions. To improve SDS, HeadSculpt [32] introduces prior-driven score distillation with landmark-guided ControlNet to improve 3D-head generation results. HeadArtist [33] develops self-score distillation for generating geometry and texture, leading to a significant performance boost.
III Preliminaries
III-A ControlNet
@@
ControlNet [34] is an innovative approach that enhances Stable Diffusion by incorporating additional conditions(canny edge, pose maps, etc.). In MaTe3D, we train two ControlNets on color mask for face and cat, respectively. The mask representation not only supports interactive editing but also enables inpainting pipeline. To achieve this, we employ a mask encoder to map the condition (e.g., mask) to the same dimensions as the latent code in Stable Diffusion. The optimization process can be expressed as follows:
@@
III-B Score Distillation Sampling (SDS)
Score Distillation Sampling (SDS) is developed by DreamFusion [28] to distill knowledge from a pre-trained diffusion model into a differentiable 3D representation generator, such as NeRF [15] or DMTet [35]. Consider a 3D representation generator $g$ is parameterized by $\theta$, we can obtain the rendered image $I=g(\theta)$. Then, given the noisy image $z_{t}$, text embedding $y$ and noise timestep $t$, a random noise $\epsilon$ is added to latent code $z$ encoded from image $I$. The pre-trained diffusion model $\epsilon_{\phi}$ is then adopted to predict the added noise from the noisy image. The SDS provides gradient to update the generator and can be represented as follows:
@@ IV Methods
Overview.   We outline the proposed MaTe3D pipeline in Fig. 1, which consists of two key parts: the proposed generator in the first step (Fig. 1(a)) and the inference-optimized editing (Fig. 1(b)) in the second step. In  section IV-A, we provide more details of the proposed generator, including the generator architecture, the SDF-based neural rendering and the training objectives. In  section IV-B, using the well-trained generator from the first step, we introduce to fuse a frozen generator and a learnable generator, and a Conditional Distillation on Geometry and Texture (CDGT) method to update the fused generator.
IV-A MaTe3D Generator
@@
IV-A1 Dual Tri-planes Generation
@@ IV-A2 Neural Rendering
Inspired by both StyleSDF [6] and VolSDF [36], we transform the signed distance values into density values for volume rendering. This transformation is formulated as follows:
@@
where $\sigma_{g}$ is the density value of the global face, and the $\sigma_{li}$ ($i=1,...,n$) refer to local facial components. The learnable hyper-parameter $\alpha$ controls the density value tightness around the surface boundary [6]. $\alpha$ is shared among both global SDF values and the decomposed local SDF values.
@@
Inspired by scene representation in [37], we directly convert SDF into semantic probabilities in generative radiance field:
@@ IV-A3 Optimization
We introduce three regularization losses: Eikonal loss ($\mathcal{L}_{eik}$) to ensure the physical valid of SDFs [6], minimal surface loss ($\mathcal{L}_{surf}$) to prevent the SDFs from modeling spurious and non-visible surfaces, and density regularization loss ($\mathcal{L}_{reg}$) to regularize the density convert from SDF to prevent ”seam” artifacts.
@@
IV-B Inference-optimized Editing
@@
Generator Fusion   Generator Fusion uses the camera pose $v$ and latent code $w$ to generate pairs of texture tri-planes and shape tri-planes. It then fuses the tri-planes by incorporating 3D masks derived from semantic probabilities. In Fig. 1(b), we illustrate the Generator Fusion of inference-optimized editing phase in MaTe3D, which comprises a frozen generator $G_{frz}$ and a learnable generator $G_{opt}$. Both $G_{frz}$ and $G_{opt}$ are initialized based on the well-trained generator from  section IV-A. We utilize $G_{frz}$ and $G_{opt}$ to derive 3D masks that facilitate feature fusion across tri-planes from both generators, allowing us to synthesize edited images $I$.
@@
IV-B1 Condition Updating Scheme
The original implementation of SDS loss primarily focuses on texture, lacking the ability to control shape [32]. Some methods enhance diffusion by incorporating landmarks obtained from an off-shelf detector for better control. However, these approaches still face challenges with unstable texture due to sparse conditions and limited 3D consistency. To tackle this issue, we introduce Condition Updating Scheme. This technology regards the generated mask $m$ as a condition for ControlNet, establishing a connection between the generator and diffusion prior. We iteratively update the condition of ControlNet during editing and integrate it into multi-view consistency to ensure stable texture. This design progressively integrates diffusion priors conditioned by 3D-aware information into the 3D content. Despite initially computing incorrect gradients during optimization, the interplay between front-view supervision and Condition Updating Scheme guarantees that results converge to maintain texture consistency and align with the modified mask.
IV-B2 Gradient Combination
While editing a neural radiance field, maintaining convincing geometry deformation is as essential as changing texture. However, we observe that the quality of geometry degrades when editing certain complex attributes (e.g., hat in Fig. 10). This degradation is attributed to a mismatch between texture and geometry caused by the absence of 3D supervision. To address this issue, we align geometry and texture by computing gradients on both the image $I$ and normal map $N$. In the approach, we introduce a novel random blending strategy to combine gradients. Specifically, we encode the image $I$ and normal map $N$ to obtain their latent representations $z^{I}$ and $z^{N}$. We then use a sampled value $\mu$ to blend texture and geometry: $\hat{z}=\mu z^{I}+(1-\mu)z^{N}$. In contrast that TADA [38] fuses texture and geometry using a constant value, our method employs randomly sampled values $\mu$ ranging from 00 to $1$. By incorporating randomness into the fusion process, random blending enhances the robustness of editing process by preventing overfitting on specific patterns and improving its capability to handle discrete attributes.
@@ IV-B2 Gradient Combination
V Experiments
V-A Implementation Details
V-A1 Setting of Generator Training
We train our generator on FFHQ [1] and proposed CatMask-HQ.We adopt sphere initialization from StyleSDF [6]. The training is performed on $4$ NVIDIA A100 GPUs ($80$G) using batch size of $8$ for $50,000$ steps. We use Adam optimizer [39] and set the learning rates for the generator and dual discriminator to $0.0025$ and $0.002$, respectively. For losses, we assign $\lambda_{adv}=1$, $\lambda_{sdf}=0.1$, $\lambda_{\sigma}=0.001$, $\lambda_{eik}=0.05$, $\lambda_{surf}=0.1$, and $\lambda_{reg}=1$. The weights $\lambda_{adv}$, $\lambda_{eik}$, $\lambda_{surf}$ and $\lambda_{reg}$ are defined the same as those in StyleSDF [6]. We define both $\lambda_{sdf}$ and $\lambda_{\sigma}$ by tentative experiments.
@@ V-A1 Setting of Generator Training
V-A2 Setting of Editing
During editing phase, we optimize learnable generator $2,000$ iterations for each target mask and input text prompt, which takes around $20$ minutes on a single NVIDIA A100 GPU ($80$G). We use Adam [39] as the optimizer with learning rate of $33e-$ 3. We determine the values of $\lambda_{CDGT}$, $\lambda_{ID}$, and $\lambda_{seg}$ as $0.001$, $0.5$, and $1$ through preliminary experiments.

### v4

@@ Abstract
3D-aware portrait editing has a wide range of applications in multiple fields. However, current approaches are limited due that they can only perform mask-guided or text-based editing. Even by fusing the two procedures into a model, the editing quality and stability cannot be ensured. To address this limitation, we propose MaTe3D: mask-guided text-based 3D-aware portrait editing. In this framework, first, we introduce a new SDF-based 3D generator which learns local and global representations with proposed SDF and density consistency losses. This enhances masked-based editing in local areas; second, we present a novel distillation strategy: Conditional Distillation on Geometry and Texture (CDGT). Compared to exiting distillation strategies, it mitigates visual ambiguity and avoids mismatch between texture and geometry, thereby producing stable texture and convincing geometry while editing. Additionally, we create the CatMask-HQ dataset, a large-scale high-resolution cat face annotation for exploration of model generalization and expansion. We perform expensive experiments on both the FFHQ and CatMask-HQ datasets to demonstrate the editing quality and stability of the proposed method. Our method faithfully generates a 3D-aware edited face image based on a modified mask and a text prompt. Our code and models will be publicly at link.
Keywords:
@@
Mask-guided Editing Text-guided Editing Diffusion Model Score Distillation Sampling
1 Introduction
@@
Leveraging the advancements of StyleGAN series stylegan stylegan2 stylegan3 zhou2022sd , EG3D eg3d and related works like stylesdf stylenerf cips3d meshwgan have successfully created high-quality 3D scenes with consistent views and rich details. Recent 3D-aware image synthesis methods fenerf ide3d lenerf aim to enhance view-consistent image editing capabilities by incorporating additional information like semantic maps. This caters to a range of user needs, including 3D avatar creation, personalized services, and industrial design.
@@
Existing NeRF-based generation methods with editability can be categorized into two classes: mask-guided fenerf ,ide3d and text-guided clipnerf lenerf . Both methods rely on a common pipeline: (1) training a 3D-aware generator; (2) exploring various editing methods based on the well-trained generator. Mask-guided methods aim to facilitate user-friendly editing through hand-drawn sketches or semantic label interfaces. For example, IDE-3D ide3d enables both high-quality and real-time local face control by manipulating a semantic map. However, current mask-guided approaches primarily emphasize shape manipulation and lack texture control capabilities. On the other hand, text-guided methods strive to produce more expressive results based on textual descriptions by distilling knowledge from large models. LENeRF lenerf utilizes CLIP clip for precise and localized manipulation using text inputs. Nevertheless, existing text-based techniques encounter challenges in accurately controlling shape to achieve desired results(Fig. 6). In this paper, we focus on using both masked-guided and text-guided 3D image editing with high quality and stability. Masks allow for precise spatial modifications, such as removing or altering specific parts of the scene, while text editing allows for semantic changes based on natural language descriptions. Utilizing both mask and text enhances interactivity and control by enabling users to interact with and modify 3D scenes in more intuitive ways.
@@
2 Related Works
2.1 3D-aware GAN based Synthesis and Editing
@@ 2 Related Works
Shocked by NeRF nerf and SIREN siren in exceeding the past method with regards to multi-view image synthesis, recent methods start to focus on 3D-aware models without multi-view supervision. GRAF graf and pi-GAN pigan are pioneers in exploring NeRF-based 3D-aware synthesis and render view-consistent results. StyleNeRF stylenerf and CIPS-3D cips3d adopt progressively upsampling to synthesize high-quality 3D-aware images. StyleSDF stylesdf , which merges NeRF with Signed Distance Function (SDF) deepsdf , achieves comparative performance to aforementioned methods and also produces 3D smooth complex shapes. EG3D eg3d introduces a hybrid explicit-implicit 3D representation that produces high-resolution image and brings implicitly learned mesh to a new level of meticulosity. However, none of the above methods consider fine-grained locally-controllable face editing with other conditions.
@@
To achieve fine-grained level control with mask, FENeRF fenerf pioneers local face editing using decoupled latent codes to generate corresponding facial semantics and texture in a spatially aligned 3D volume with shared geometry. IDE-3D ide3d and NeRFFaceEditing nerffaceediting allow for semantic mask guided local control of images in real time. CNeRF cnerf and 3D-SSGAN 3dssgan aim to achieve part-level editing through the compositional synthesis. For fine-grained editing with text, LENeRF lenerf achieves local manipulation via establishing 3D attention map via CLIP clip . StyleGANFusion styleganfusion and FaceDNeRF facednerf distill knowledge from diffusion model to enable manipulate Face NeRFs in zero-shot learning based on text prompts.
2.2 Diffusion Models Guided Editing
@@
In recent years, diffusion models have made great strides in the field of text-to-2D generation. Some works have begun to use diffusion model priors to edit 3D scenes through text guidance. For example, Instruct-NeRF2NeRF haque2023instruct employs Instruct-Pix2Pix instructp2p to modify NeRF content using language instructions. The approach optimizes both the underlying scene and iterative optimization, resulting in a 3D scene that conforms to the input text. Blended-NeRF gordon2023blended uses a pre-trained language image model to synthesize objects based on textual prompts and blend them into existing scenes using 3D multi-layer perceptron (MLP) models and volume blending techniques for natural appearance.
@@
Score Distillation Sampling (SDS) from DreamFusion dreamfusion provides a new way to apply diffusion models to 3D scenes. DreamEditor zhuang2023dreameditor allows controllable editing of neural fields using textual prompts through SDS. FocalDreamer li2023focaldreamer merges shapes with editable parts based on SDS for fine-grained editing in desired regions. SKED mikaeili2023sked , which edits 3D shapes represented by NeRFs, utilizes two guiding sketches without altering existing fields much, supporting object creation or modification within designated sketch regions. To improve SDS, HeadSculpt headsculpt introduces prior-driven score distillation with landmark-guided ControlNet to improve 3D-head generation results. HeadArtist headartist develops self-score distillation for generating geometry and texture, leading to a significant performance boost.
3 Preliminaries
3.1 ControlNet
@@
ControlNet controlnet is an innovative approach that enhances Stable Diffusion by incorporating additional conditions(canny edge, pose maps, etc.). In MaTe3D, we train two ControlNets on color mask for face and cat, respectively. The mask representation not only supports interactive editing but also enables inpainting pipeline. To achieve this, we employ a mask encoder to map the condition (e.g., mask) to the same dimensions as the latent code in Stable Diffusion. The optimization process can be expressed as follows:
@@
3.2 Score Distillation Sampling (SDS)
Score Distillation Sampling (SDS) is developed by DreamFusion dreamfusion to distill knowledge from a pre-trained diffusion model into a differentiable 3D representation generator, such as NeRF nerf or DMTet dmtet . Consider a 3D representation generator $g$ is parameterized by $\theta$, we can obtain the rendered image $I=g(\theta)$. Then, given the noisy image $z_{t}$, text embedding $y$ and noise timestep $t$, a random noise $\epsilon$ is added to latent code $z$ encoded from image $I$. The pre-trained diffusion model $\epsilon_{\phi}$ is then adopted to predict the added noise from the noisy image. The SDS provides gradient to update the generator and can be represented as follows:
@@ 4 Methods
Overview.   We outline the proposed MaTe3D pipeline in Fig. 1, which consists of two key parts: the proposed generator in the first step (Fig. 1(a)) and the inference-optimized editing (Fig. 1(b)) in the second step. In Sec. 4.1, we provide more details of the proposed generator, including the generator architecture, the SDF-based neural rendering and the training objectives. In Sec. 4.2, using the well-trained generator from the first step, we introduce to fuse a frozen generator and a learnable generator, and a Conditional Distillation on Geometry and Texture (CDGT) method to update the fused generator.
4.1 MaTe3D Generator
@@
4.1.1 Dual Tri-planes Generation
@@ 4.1.2 Neural Rendering
Inspired by both StyleSDF stylesdf and VolSDF volsdf , we transform the signed distance values into density values for volume rendering. This transformation is formulated as follows:
@@
where $\sigma_{g}$ is the density value of the global face, and the $\sigma_{li}$ ($i=1,...,n$) refer to local facial components. The learnable hyper-parameter $\alpha$ controls the density value tightness around the surface boundary stylesdf . $\alpha$ is shared among both global SDF values and the decomposed local SDF values.
@@
Inspired by scene representation in objectsdf , we directly convert SDF into semantic probabilities in generative radiance field:
@@ 4.1.3 Optimization
We introduce three regularization losses: Eikonal loss ($\mathcal{L}_{eik}$) to ensure the physical valid of SDFs stylesdf , minimal surface loss ($\mathcal{L}_{surf}$) to prevent the SDFs from modeling spurious and non-visible surfaces, and density regularization loss ($\mathcal{L}_{reg}$) to regularize the density convert from SDF to prevent ”seam” artifacts.
@@
4.2 Inference-optimized Editing
@@
Generator Fusion   Generator Fusion uses the camera pose $v$ and latent code $w$ to generate pairs of texture tri-planes and shape tri-planes. It then fuses the tri-planes by incorporating 3D masks derived from semantic probabilities. In Fig. 1(b), we illustrate the Generator Fusion of inference-optimized editing phase in MaTe3D, which comprises a frozen generator $G_{frz}$ and a learnable generator $G_{opt}$. Both $G_{frz}$ and $G_{opt}$ are initialized based on the well-trained generator from Sec. 4.1. We utilize $G_{frz}$ and $G_{opt}$ to derive 3D masks that facilitate feature fusion across tri-planes from both generators, allowing us to synthesize edited images $I$.
@@
4.2.1 Condition Updating Scheme
The original implementation of SDS loss primarily focuses on texture, lacking the ability to control shape headsculpt . Some methods enhance diffusion by incorporating landmarks obtained from an off-shelf detector for better control. However, these approaches still face challenges with unstable texture due to sparse conditions and limited 3D consistency. To tackle this issue, we introduce Condition Updating Scheme. This technology regards the generated mask $m$ as a condition for ControlNet, establishing a connection between the generator and diffusion prior. We iteratively update the condition of ControlNet during editing and integrate it into multi-view consistency to ensure stable texture. This design progressively integrates diffusion priors conditioned by 3D-aware information into the 3D content. Despite initially computing incorrect gradients during optimization, the interplay between front-view supervision and Condition Updating Scheme guarantees that results converge to maintain texture consistency and align with the modified mask.
4.2.2 Gradient Combination
While editing a neural radiance field, maintaining convincing geometry deformation is as essential as changing texture. However, we observe that the quality of geometry degrades when editing certain complex attributes (e.g., hat in Fig. 10). This degradation is attributed to a mismatch between texture and geometry caused by the absence of 3D supervision. To address this issue, we align geometry and texture by computing gradients on both the image $I$ and normal map $N$. In the approach, we introduce a novel random blending strategy to combine gradients. Specifically, we encode the image $I$ and normal map $N$ to obtain their latent representations $z^{I}$ and $z^{N}$. We then use a sampled value $\mu$ to blend texture and geometry: $\hat{z}=\mu z^{I}+(1-\mu)z^{N}$. In contrast that TADA tada fuses texture and geometry using a constant value, our method employs randomly sampled values $\mu$ ranging from 00 to $1$. By incorporating randomness into the fusion process, random blending enhances the robustness of editing process by preventing overfitting on specific patterns and improving its capability to handle discrete attributes.
@@ 4.2.2 Gradient Combination
5 Experiments
5.1 Implementation Details
5.1.1 Setting of Generator Training
We train our generator on FFHQ stylegan and proposed CatMask-HQ.We adopt sphere initialization from StyleSDF stylesdf . The training is performed on $4$ NVIDIA A100 GPUs ($80$G) using batch size of $8$ for $50,000$ steps. We use Adam optimizer adam and set the learning rates for the generator and dual discriminator to $0.0025$ and $0.002$, respectively. For losses, we assign $\lambda_{adv}=1$, $\lambda_{sdf}=0.1$, $\lambda_{\sigma}=0.001$, $\lambda_{eik}=0.05$, $\lambda_{surf}=0.1$, and $\lambda_{reg}=1$. The weights $\lambda_{adv}$, $\lambda_{eik}$, $\lambda_{surf}$ and $\lambda_{reg}$ are defined the same as those in StyleSDF stylesdf . We define both $\lambda_{sdf}$ and $\lambda_{\sigma}$ by tentative experiments.
@@ 5.1.1 Setting of Generator Training
5.1.2 Setting of Editing
During editing phase, we optimize learnable generator $2,000$ iterations for each target mask and input text prompt, which takes around $20$ minutes on a single NVIDIA A100 GPU ($80$G). We use Adam adam as the optimizer with learning rate of $33e-$ 3. We determine the values of $\lambda_{CDGT}$, $\lambda_{ID}$, and $\lambda_{seg}$ as $0.001$, $0.5$, and $1$ through preliminary experiments.